OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 3, 7, the square root of 50 A. yes B. no

OpenStudy (anonymous):

a^2 + b^2 = c^2 or 3^2 + 7^2 = sqrt(50)^2

OpenStudy (anonymous):

3^2 + 7^2 = 9 + 49 = 58 sqrt(58) does not equal to square root of 50 so it's no

OpenStudy (anonymous):

which means 9 +49 = 50 which is not true. The answer is no.